### v2

# Trust-Preserved Human-Robot Shared Autonomy enabled by Bayesian Relational Event Modeling

Abstract
Shared autonomy functions as a flexible framework that empowers robots to operate across a spectrum of autonomy levels, allowing for efficient task execution with minimal human oversight. However, humans might be intimidated by the autonomous decision-making capabilities of robots due to perceived risks and a lack of trust. This paper proposed a trust-preserved shared autonomy strategy that allows robots to seamlessly adjust their autonomy level, striving to optimize team performance and enhance their acceptance among human collaborators. By enhancing the relational event modeling framework with Bayesian learning techniques, this paper enables dynamic inference of human trust based solely on time-stamped relational events communicated within human-robot teams. Adopting a longitudinal perspective on trust development and calibration in human-robot teams, the proposed trust-preserved shared autonomy strategy warrants robots to actively establish, maintain, and repair human trust, rather than merely passively adapting to it. We validate the effectiveness of the proposed approach through a user study on a human-robot collaborative search and rescue scenario. The objective and subjective evaluations demonstrate its merits on both task execution and user acceptability over the baseline approach that does not consider the preservation of trust.
I Introduction
As robots continue to find applications in a wide range of real-world scenarios, the next wave of technological progress in the field of robotics appears to be centered around their ability to interact effectively with complex environments and autonomously execute tasks with minimal human supervision [1, 2, 3]. Although maintaining human global control over autonomous systems contributes to increasing the initiative and awareness of dynamic, complex, and interactive environments, the ability of robots to make unaided decisions also grows in significance [4, 5, 6].
On the one hand, the efficiency of the human-robot teams can be compromised due to frequent human intervention [7, 8, 9]. From the human perspective, overseeing and instructing robots, particularly in multi-robot teams where a large group of robots require simultaneous management, can demand a significant commitment of effort. For robots, continuous interruptions from human can cause disturbances that make their operation destabilized, dependent, and slow [10], especially in the presence of communication delays.
On the other hand, the fundamental expectation behind human-robot teaming is that the abilities of robots and humans complement each other, resulting in enhanced performance compared to their individual efforts [11, 12, 13]. Robots can often offer an information advantage by accessing data that are unavailable to human team members [10]. For example, consider a search and rescue (SAR) scenario as illustrated in Fig. 1. The human responder might initially focus on rescuing injured individuals. At the same time, the robot identifies a gas leakage that requires immediate attention to prevent a potential explosion endangering everyone within the scene. In such a situation, the robot is challenged to weigh the importance of two conflicting goals: adhering to the human’s instructions to continue rescuing the victim versus prioritizing putting down the gas leakage. Hence in the best interests of the team as a whole, the robots are expected to make judicious and rational assessments to determine when to adhere to or disregard human directives when conflicts happen.
Figure 1: A search and rescue scenario with human-robot team. Robots may acquire information unavailable to human team members, and make strictly rational decisions.
Shared Autonomy (SA). When teaming up with robots, human’s expectations of robots now have gone beyond merely an “extended hand” to alleviate physical tasks, but an “extended brain” that can support, inform, and even rectify us when necessary [14, 15, 16]. The degree to which the robots autonomously choose to follow human instructions in the light of other objectives can be characterized by their autonomy level [7]. Shared autonomy provides an opportunity to establish a flexible autonomy framework that grants robots the capacity to function across varying autonomy levels [3]. There have been various autonomy adaptation strategies proposed utilizing information extracted from both the human (e.g. intentions, muscle activity, skills) and the environment (e.g. task, constraints) [17, 18, 19, 20, 21, 22].
However, most of those approaches ignore the fact that the autonomous decision-making capabilities of robotic systems can sometimes be unsettling for humans due to various factors, including perceived risks and a lack of trust [23]. Therefore, to achieve more fluent human-robot teaming, concerns regarding the acceptability of non-human entities making decisions autonomously on behalf of human team members also need to be addressed. From a longitudinal perspective on trust development and calibration in human-robot teams [24], the robot’s autonomy level should be well aligned with the human’s trust level to avoid autonomy surprises. Consequently, this paper paid special attention to the maintenance of human trust during the adaptation of the robot’s autonomy level.
Trust Inference. The main challenge that obstructs successful autonomy-trust coordination is the inherent difficulty of quantifying human trust. Numerous researchers have approached this challenge from diverse viewpoints, for example, physiological measures (e.g. skin conductance, eye-tracking), neuroimaging, behavioral observations, and self-reported trust ratings [25, 26, 27]. Those approaches function effectively in the scenarios of proximate human-robot interaction where humans and robots share the same workspace, allowing for ample human observations.
However, the majority of those methods lose their throne in more general real-world scenarios of human-robot teaming where humans and robots are often separated spatially and temporally. For example, in the SAR scenarios, disaster robots are separated from their remote operators in both space and time. In those cases, the only available interaction information is a flow of commands communicated between humans and robots, where none of these events by itself characterizes the team relationships at any point in time. To tackle this challenge, this paper proposed an innovative approach to dynamically infer the underlying drivers of interaction dynamics (e.g. trust) based merely on time-stamped relational events communicated within human-robot teams (e.g. commands).
The contributions of this paper are summarized as:
Trust Inference from Remote Interactions. We propose a novel approach to dynamically infer human trust from remote human-robot interactions by enhancing the relational event modeling (REM) framework [28, 29, 30, 31, 32, 33] with Bayesian learning techniques.
Trust-preserved Shared Autonomy. We devise a trust-preserved shared autonomy strategy that allows the robots to seamlessly adjust their autonomy level while actively establishing, maintaining, and repairing human trust.
User Study on Collaborative SAR. We validate the effectiveness of the proposed method by performing a user study on a human-robot collaborative SAR scenario.
II Problem Formulation
We are inspired by a human-robot collaborative SAR scenario where the human and multiple robots jointly perform the search and evacuation of victims from an incident area [34], as illustrated in Fig. 2.
Figure 2: A human-robot collaborative SAR scenario where the human and multiple robots jointly perform the search and evacuation of victims from an incident area.
Environment: There are several buildings that have been hit by an earthquake, and whose entrances may be blocked by fire. Each building may contain one or more victims, who may be injured or uninjured. One of the buildings is undergoing gas leakage inside, and if the gas density achieves a certain level, it will lead to an explosion.
Task: The team needs to localize all victims in the buildings and assess their conditions. The injured victims should be treated and escorted to the shelter.
Sub-task sequence dependencies: The victims blocked by fire cannot be accessed until the fire blocking the entrance is put down. The injured victims must be treated before being transmitted to the shelter.
Team member capabilities: The capabilities of the robots and humans are complementary. Only robots can sense and estimate the location and density of the gas leakage. Furthermore, they can put down fires and carry injured victims, but humans cannot. In contrast, humans can treat injured victims, which cannot be done by robots. However, both humans and robots can search and assess the victims, as well as shut down the gas leakage. In this way, the task can only be achieved by the collaboration of different types of team members.
Human-robot communications: Humans and robots can communicate through commands using template sentences.
The human has the option to either acquire the robots’ status or give instructions to the robots.
The human can inquire about the status of robots by sending a command to a certain robot asking “What are you doing?” As a response, the inquired robot will provide the current sub-task that it is performing, for example, “I’m going to Building X to put down the fire,” or “I’m escorting the injured victim to the shelter.”
The human can also instruct a certain robot by telling that robot to “Go to Building X.” If the instruction from the human conflicts with the robot’s current sub-task, it can either obey the human’s instruction regardless of its own mission by responding with “Sure, I am going there.”; or stick to its current mission and convey it to the human “Sorry, but I have to prioritize going to Building X.”
Furthermore, the robots can also share their understanding of the environment with the human (e.g., “Instruction received. However, I have the following additional environmental information that may be useful for you to refine the decisions. Please resend your decision.”); and provide an explanation of their decisions to the human (e.g., “My priority has been adjusted due to an observed dangerous condition in Building X.”).
Problem Statement: Grounded on the toy example scenario described above, the objective of this paper is to develop a trust-preserved shared autonomy framework that can:
  • (i)
    Dynamically infer the human’s trust in the robots based on merely time-stamped commands communicated between the human and robots;
  • (ii)
    Adaptively coordinate the robot’s autonomy level with the human’s trust level, which facilitates both the completion of the task and the acceptability of the human.
III Bayesian Relational Event Modelling
III-A Relational Event Modelling
Interactions among individuals can be represented as a sequence of relational event, which is defined as “a discrete event generated by a social actor and directed toward one or more targets” [29]. Formally, a relational event is represented by a tuple $e=(i_{e},j_{e},k_{e},w_{e},t_{e})$ containing the sender $i_{e}$, receiver $j_{e}$, type $k_{e}$, weight $w_{e}$, and time $t_{e}$ of an event. The full sequence of $m$ events formulates the history event set $E=\{e_{1},\cdots,e_{m}\}$. Let $A=\{1,\cdots,n\}$ be the set of all $n$ actors, and $D\subseteq A\times A$ the set of all dyads. Events can fall into a set of $K$ discrete classes. For each class $k$, define the weight function $w_{kt}$ as the accumulated weight of all past interaction events [35].
Such events constitute a dynamic network of directional ties, where the probability of an event occurring often depends on various kinds of attributes in the network. For example, individual covariates such as gender, age, or institutional role; dyadic covariates such as friendship status or trust; environmental factors including information about the physical environment, the stage in a creative process, or the status of an emergency. For simplicity of notation, we combine all those attribute information into a set $G_{t}$.
REM is a blend of social sequence and network analysis, which provides a principled approach to investigate the underlying factors that drive the occurrence, rhythm, and speed of individuals interacting with each other over time. REM models every potential pairing event of individuals in terms of rates, where common events have high rates, and rare events have low rates. Taking from the Cox proportional hazards model [36], the rate for a sender $i$ directing an action of type $k$ to receiver $j$ at time $t$ is modeled as a log-linear function of a set of sufficient statistics $S_{p}(i,j,k,G_{t})$ and their corresponding intensity parameters $\theta_{p}$:
$\lambda_{ijk}(t|G_{t},\boldsymbol{\theta})=\lambda_{0}(t)\exp{\left(\sum_{p=1}^{P}\theta_{p}S_{p}(i,j,k,G_{t})\right)},$
where $\boldsymbol{\theta}=[\theta_{1},\cdots,\theta_{P}]^{\intercal}$, and the parameter $\lambda_{0}(t)$ controls the relationship between the rate and time. The simplest case is to assume that this value is constant, which corresponds to the exponential model.
The likelihood function for the full event sequence and timing can be explicitly computed as the probability of each event, multiplied by the probability of the time between events, conditional on the entire realized history. Leaving out the time-dependent parameter $\lambda_{0}(t)$, the partial likelihood function is given by
$\begin{split}&f(E;G_{t},\boldsymbol{\theta})=\prod_{e\in E}\lambda_{i_{e}j_{e}k_{e}}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})\times\\ &\exp{\left(-\Delta t_{e}\sum_{k=1,\cdots,K}\sum_{(i,j)\in D}\lambda_{ijk}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})\right)},\end{split}$ (1)
where $\Delta t_{e}=t_{e}-t_{e-1}$ is the time elapsed between events. This expression represents the likelihood of every event in the sequence, along with the conditional likelihood that no other events occurred in the period between observed events.
If the relational event data is simply ordinal, i.e. there is no specific timing data available, the likelihood function can be reduced to a product of multinomial probabilities derived from each event rate
$f(E;G_{t},\boldsymbol{\theta})=\prod_{e\in E}\frac{\lambda_{i_{e}j_{e}k_{e}}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})}{\sum_{k=1,\cdots,K}\sum_{(i,j)\in D}\lambda_{ijk}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})}.$
III-B Sufficient Statistics
The greatest utility of REM lies in the wide range of sufficient statistics that can be derived from event sequences. These measures are numerical representations of specific interaction patterns, similar to those encoded in exponential random graph models (ERGMs) or stochastic actor-oriented models [37]. That is, like how conceptual models are translated into statistical models, sufficient statistics are the operationalizations of model parameters including both endogenous and exogenous factors.
The endogenous factors summarize how the accumulation and sequencing of past relational events can influence the likelihood of the next relational event. For instance, inertia reflects the degree to which group members’ past contacts tend to be their future contacts; Reciprocity examines the likelihood of group member B sending a message to A if B had recently received a message from A.
Exogenous factors refer to characteristics outside the relational event history, including individual attributes, relational attributes, as well as environmental contextual information beyond the scope of the interaction system. For example, sometimes the likelihood of a group interaction occurring is due to some individual attributes of either the sender or receiver. An extroverted individual might send messages at a higher rate than somebody who is more introverted. Relational attributes refer to different kinds of ties individuals might have with one another like affinity (e.g., friendship, trust), flow (e.g., other forms of sending messages), representational (e.g., endorsements), and semantic (e.g., shared interpretations) ties [38]. Environmental information may encompass more readily measurable entities like the nature of the task, restrictions on communication channels, or availability of resources.
Fig. 3 describes some sufficient statistics used to model the group communications for the collaborative SAR scenario investigated in this paper. For endogenous factors, we include inertia and reciprocity. For exogenous factors, we include the human’s individual attribute of command and the relational attribute of trust between the human and the robots. Those sufficient statistics are by no means exhaustive, and various other structures may be encoded to capture specific behavioral patterns if needed.
Figure 3: Examples of sufficient statistics for the collaborative SAR scenario. The solid arrows represent past historical events, and the dashed arrows represent the future events to be predicted. $\mu(i,j,k,t)$ measures the frequency of event $(i,j,k)$ in the past event history before $t$.
III-C Bayesian Relational Event Modelling
To dynamically infer the evolution of human trust, we propose to advance REM into an adaptive framework taking advantage of Bayesian learning. This framework is composed of two stages: 1) offline learning for team properties grounding, and 2) online learning for dynamic attributes inference.
Team properties grounding. The sign and magnitude of each element in the coefficient parameter $\boldsymbol{\theta}$ determines how influential a particular network effect is regarding the generation of relational events across team members. Therefore, the first stage aims to estimate the coefficient parameter $\boldsymbol{\theta}$ to ground the specific properties of the team to be investigated. In this stage, a period of interaction history $E_{0}$ within the team is collected, as well as the initial attribute information $G_{0}$ that can be obtained by a post-process survey. Then the coefficient estimates for relational event models – ordinal or temporal – are derived from the solution of the maximum likelihood problem:
$\boldsymbol{\theta}^{\ast}:=\arg\max_{\boldsymbol{\theta}}\log\left(f(E_{0};G_{0},\boldsymbol{\theta})\right).$
Common optimization techniques such as Newton-Rhapson are well-suited to solving this maximization problem.
Dynamic attributes inference. Based on the grounded properties $\boldsymbol{\theta}^{\ast}$ for a specific team, the next stage is to deploy a receding horizon approach to dynamically infer the time-varying attributes $G_{t}$ in the team (e.g., trust), as illustrated in Fig. 4. We assume that within each observation interval $T_{k}$ between two discrete observation time instant $t_{k-1}$ and $t_{k}$, the attributes $G_{t}$ remain constant. And the observed relational events history within the observation interval $T_{k}$ is recorded as $E_{k}$. Define $\pi_{k}(G_{t})=P(G_{t}|E_{1:k},\boldsymbol{\theta}^{\ast})$ as the posterior distribution of $G_{t}$ after $k$-th observation intervals. According to the Bayes’ Theorem,
$\begin{split}\pi_{k}(G_{t})&=P(G_{t}|E_{1:k},\boldsymbol{\theta}^{\ast})\\ &=\frac{P(E_{k}|G_{t},E_{1:k-1},\boldsymbol{\theta}^{\ast})P(G_{t}|E_{1:k-1},\boldsymbol{\theta}^{\ast})}{\sum_{G_{t}}P(E_{k}|G_{t},E_{1:k-1},\boldsymbol{\theta}^{\ast})P(G_{t}|E_{1:k-1},\boldsymbol{\theta}^{\ast})}\\ &=\frac{P(E_{k}|G_{t},\boldsymbol{\theta}^{\ast})\pi_{k-1}(G_{t})}{\sum_{G_{t}}P(E_{k}|G_{t},\boldsymbol{\theta}^{\ast})\pi_{k-1}(G_{t})}\\ &=\frac{f(E_{k};G_{t},\boldsymbol{\theta}^{\ast})\pi_{k-1}(G_{t})}{\sum_{G_{t}}f(E_{k};G_{t},\boldsymbol{\theta}^{\ast})\pi_{k-1}(G_{t})},\end{split}$
where the likelihood function $f(E_{k};G_{t},\boldsymbol{\theta}^{\ast})$ can be calculated according to (1), and the prior distribution $\pi_{0}(G_{t})$ offers the opportunity to take the initial attribute information $G_{0}$ collected in the previous stage into consideration.
Figure 4: Bayesian REM framework. A moving window technique and sequential Bayesian update rule are utilized to update the estimation of team attributes online based on newly observed relational events.
IV Trust-Preserved Shared Autonomy
IV-A Quantification of Trust and Autonomy Levels
Following [39], we define trust as “the willingness of a party to be vulnerable to the actions of another party based on the expectation that the other will perform a particular action important to the trustor, irrespective of the ability to monitor or control that other party”. By including the human trust $\beta$ into the time-varying attributes $G_{t}$, we can dynamically infer the evolution of trust based on Bayesian REM. Superior to some point estimation approaches, e.g., maximum likelihood estimation (MLE), the utilization of Bayesian learning enables us to maintain the complete statistic description of trust through a probability distribution $\pi(\beta)$, which offers greater flexibility in quantifying trust level tailored to specific scenarios. For example, the trust level $L_{\beta}$ in general can be determined by taking expectation of the estimated trust distribution: $L_{\beta}=\mathbb{E}_{\pi(\beta)}\beta$. However, in some safety-critical scenarios, we may prefer to take a risk-averse perspective with certain risk measures: $L_{\beta}=\mathcal{R}_{\pi(\beta)}\beta$, such as mean-variance, Value at Risk (VaR), and Conditional Value at Risk (CVaR) (cf. [40]).
Level of robot autonomy, ranging from teleoperation to fully autonomous systems, defines the extent to which a system can carry out its own processes and operations without external control. There have been different criteria of assessing the autonomy level, for example, the allocation between functions, or the amount of time that a person can neglect the robot [7]. In the context of this paper, we quantify the autonomy level $L_{\alpha}$ of the robot as $L_{\alpha}=1-\alpha$, where $\alpha$ is the probability that the robot follows the human’s instruction when contradicting its own reasoning.
IV-B Coordination between Trust and Autonomy Levels
After formally quantifying the human trust level and robot autonomy level, the immediate question that arises is: How to effectively adjust the robot autonomy level according to the evolving human trust level? In general, the more trust humans place on robots, the higher level of autonomy they tend to accept. Therefore, an intuitive adaptive strategy is to passively choose the autonomy level that best matches the estimated trust level: $L_{\alpha}\propto L_{\beta}$, as illustrated in Fig. 5.
However, this reactive adjustment may impede the effectiveness and efficiency of the human-robot team and fail to take full advantage of the robot’s capabilities. For example, when humans maintain a lower level of trust in the robots, disuse or micromanagement may happen, which can result in suboptimal solutions to both task completion and increased workload for individual members.
Figure 5: The coordination between the robot’s autonomy level and the human’s trust level.
In real-world human-robot teaming scenarios, the collaboration between humans and robots is a long-term dynamic process, where all team members continuously engage in the process of establishing and calibrating trust among each other. Therefore, it is expected for robots to deploy a trust-preserve strategy to actively establish, maintain, and repair human trust, rather than merely passively adapting to it.
IV-C Trust-preserved Shared Autonomy
Methods for trust repair
Description
Apology
Admission of fault and statement of remorse
Denial
Recognition of error and assertion that they are not at fault
Control
Statement of control/awareness of situation
Convey uncertainty directly
Conveying uncertainty directly by providing machine confidence indicators
Show critical states
Enhance an operator’s mental state by showing only the critical states, situations that require immediate action, of an algorithm’s policy
TABLE I: Typical trust repair strategies suitable for remote human-robot collaboration scenarios.
Figure 6: The punctuated equilibrium model of team dynamics, where the inflection point indicates a trust violation and the coming of revolutionary periods, during which some trust repair strategies should be deployed to re-establish the trust between humans and robots.
To devise a trust-preserved shared autonomy strategy, we resort to an analysis of the temporal dynamics about trust formation and maintenance within a broader perspective of longitudinal teaming. Punctuated Equilibrium Model (PEM) [41], which describes the pattern of evolution in evolutionary biology and paleontology, reveals that the dynamics of groups often involve long periods of relative stability (equilibrium) punctuated by relatively short bursts of rapid change (punctuation), as illustrated in Fig. 6. According to the paradigm of PEM, the team structure is considered stable and experiences limited change over the evolutionary periods, representing equilibrium. These periods are interrupted by brief but significant episodes of revolutionary periods, where the team structure may disassemble, reconfigure, and experience qualitative, metamorphic change. The transition from an evolutionary period into a revolutionary period can be characterized by the inflection point of the team dynamics, which is a critical moment or juncture when a significant change or turning point occurs in the team structure.
In the context of human-robot teaming, evolutionary periods represent the periods when humans and robots have established and maintained a relatively stable trust and relationship. Within those periods, robots can simply adjust the autonomy level to match the estimated trust level $L_{\alpha}\propto L_{\beta}$. However, for revolutionary periods, which often happen at the beginning of the teaming process or when trust violation occurs, additional trust repair approaches should be deployed to establish or re-establish the trust between team members.
There are various kinds of methods to repair the trust in human-robot teams, for example, by explaining the cause and/or situational nature of a mistake, or making promises about future behavior. Following [24], we summarize several typical strategies that are suitable for remote human-robot collaboration scenarios, as demonstrated in Table. I. For instance, in the scenario of collaborative SAR, the initial trust from the human may be low due to unfamiliarity. Then during this initial revolutionary period, the robots could explain their decisions to the human and seek confirmation from the human whenever making a counter-command decision. It is also possible that the robots occasionally make decisions that appear “unethical” to the human, resulting in a loss of trust. In such cases, rather than simply switching to a lower autonomy level, the robots could state their awareness of the situation and assure the human by providing the estimated probability of a dangerous condition.
V User Study
To verify whether the proposed trust-preserved shared autonomy approach actually assists a human operator in executing tasks that require a high workload and fluent human-robot collaboration, a user study was conducted on a human-robot collaborative SAR scenario.
V-A Study Design
Participants: We recruited 16 college student subjects (3 female, average ages 26.06$\pm$4.93 years) to participate in our study. All subjects were provided informed consent prior to the experiment.
Tasks: Each participant needs to play a simulated human-robot collaborative SAR game, as described in Section II.
Experimental Setup: A simulation interface to perform the SAR task was developed using the Graphical User Interfaces (GUIs) tool in Matlab, as illustrated in Fig. 7. The center figure displays the workspace of SAR, where the buildings are represented with squares. We deploy a team that contains one human and two robots, whose locations are displayed by the blue circle “H” and green circles “R1” and “R2”, respectively. The icons on the left of the workspace demonstrate the colors used to represent the different statuses of those buildings. The right interface part displays the actions that the user can conduct. In the “Command to Robot” panel, users can inquire about about each robot’s current targeted building by choosing the corresponding robot in the pop-menu and hitting the “Status?” button. Similarly, users can also instruct a certain robot to go to a certain building by changing the pop-menus and hitting the “Go!” button. The “Human” panel is used by the users to control the human’s movement, and conduct different sub-tasks.
Figure 7: The simulation interface used by the user study to perform the SAR task. The center figure displays the workspace of SAR. The icons on the left of the workspace demonstrate the colors that are used to represent different statuses of those buildings. The right interface part displays the actions that the user can conduct.
Procedure: Each participant received an instruction document that contained an explanation of the scenario and tasks, and an instruction on how to use the interface.
The participant first played the game in the practice phase to get familiar with the task and interface. In this phase, the robots were assigned two fixed autonomy levels, one with a higher autonomy level ($L_{\alpha}=0.9$) and one with a lower autonomy level ($L_{\alpha}=0.1$). To prevent the participants from learning how to go about the entire SAR operation, the scenarios (which building was burning, which one had a gas leak, where they were located, which one had people) were randomized for each trial. Participants could try as many trials as they wish at this stage, after which they were asked to report their 10-point Likert scale on their trust in each robot. The interaction and trust data collected at this stage was utilized to perform the offline learning for team properties grounding, which captured the individual differences of the participants (e.g. if a higher trust predicts the probability of sending fewer inquiring messages).
In the next stage, a within-subject design was applied, namely, each participant completed the tasks with two conditions (Baseline SA and Trust-preserved SA) for 6 trials, with the order of conditions counterbalanced across the participants. After completing the experimental tasks, participants were administered a brief questionnaire to assess their user experiences.
Variables: For comparison, we tested two conditions with different settings of robot autonomy. In the baseline condition (Baseline SA), each robot follows a fixed autonomy level according to the reported trust at the first stage. On the contrary, the robots in another condition (Trust-preserved SA) are equipped with our proposed strategy, which are capable of dynamically inferring human trust and seamlessly coordinating their autonomy level during the operation.
V-B Measures
Objective Measures: Defining a successful task execution as completing the task without a gas explosion, we recorded each participant’s success rate over multiple trials with both conditions. For each successful task execution, performance measures were obtained by automatically logging data during task execution. The following measures were collected: task duration (shorter completion time means better performance); and the number of commands sent by the human agent (fewer commands means better performance).
Subjective Measures: We administered a 10-point Likert scale survey for each condition. The test distinguishes five dimensions of collaboration fluency: How easy it was to complete the tasks; How helpful the robots were to complete the tasks; How much trust had been put into the robots about their situation awareness; In what degree the robots acted like a member of the team; In what degree the robots’ responses helped improve the trust. We also asked each participant’s preference on which condition they would choose to work with in the future.
V-C Results
The objective and subjective data are analyzed, whose means and standard deviations are displayed in Fig. 8 and Fig. 9, respectively.
Figure 8: Means and standard deviations of objective measures. Significant enhancement of success rate and reduction of workload are demonstrated.
Figure 9: Means and standard deviations of subjective measures. There are significant improvements in the users’ experience over all five dimensions of collaboration fluency.
[…]
The reason that there is no significant difference in the task duration between the two conditions may be due to the fact that only the completion time of successful task execution was recorded. In those trials, users may occasionally deploy a near-optimal strategy of task execution, which leads to a near-optimal task completion time.
[…]
VI Conclusion and Future Work
In this paper, a trust-preserved shared autonomy strategy was proposed, which allows robots to seamlessly adjust their autonomy level, striving to optimize team performance and enhance their acceptance among human collaborators. Providing an innovative approach to model the dynamics of human trust based on Bayesian REM, this paper opens a new perspective of trust modeling in remote human-robot interaction scenarios. Future work will investigate the application of the proposed method to more general and complex human-robot teaming scenarios, for example, considering the disuse of robots or dynamically changing priorities.